# Stimulus-Informed Generalized Canonical Correlation Analysis for Group Analysis of Neural Responses to Natural Stimuli

### v2

@@ 1 Introduction
In the past few years, we have seen a surge of BCI applications that are passive, i.e., tapping into the natural behavior of a user, single-trial, i.e., not requiring repetitions of the stimulus, and operate on natural sensory stimuli, such as speech/music and natural video [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. These new BCI paradigms can then be employed in much more mainstream application domains, such as hearing aids and consumer earphones [14, 3, 15, 4, 16], educational sciences [5, 11, 17], neuromarketing [6], or virtual reality environments [18]. Many of these applications require decoding stimulus-following neural responses, for example, to quantify levels of absolute [8, 10, 19] or selective attention [14, 3] to a particular auditory or visual stimulus. The temporal dynamics of the stimulus then result in a so-called stimulus-following neural response, which can be decoded from different neurorecording modalities such as electroencephalography (EEG), magnetoencephalography (MEG), or electrocorticography (ECoG). The EEG modality is particularly interesting because it is non-invasive, low cost and highly mobile [1].
However, decoding neural responses to natural stimuli is much more challenging from a signal processing perspective, given that they are much more unpredictable (resulting from the nature of the stimulus) and suffer from a very low SNR, as the targeted stimulus-following neural responses are buried under all kinds of non-neural and neural noise. Given the single-trial paradigm, averaging the responses across multiple trials is not an option anymore to deal with this extremely low SNR. Therefore, much more advanced, data-driven signal processing algorithms are required and being developed to enhance the targeted stimulus-following activity and suppress all other noise [20, 3, 21]. In this paper, we focus on the group analysis of such stimulus-following responses, i.e., we assume a set of synchronized neural responses to the same natural stimulus is available, e.g., from a group of subjects all attending to the same stimulus. The group analysis of the stimulus-following neural responses can then be a goal in itself, e.g., to decode a notion of group attention to the stimulus, as employed in educational neuroscience [5] or neuromarketing [6]. Alternatively, the group information can be leveraged to assist the decoding of stimulus-following neural responses on an individual level, using it as, e.g., a preprocessing technique to a priori improve the SNR [22, 23].
@@ 3 SI-GCCA/SI-corrCA
We now propose the stimulus-informed GCCA (SI-GCCA) technique for the group analysis of stimulus-following neural responses, in which we include the stimulus as side information, as visualized in Figure 1. The implicit assumption in SI-GCCA is, therefore, that the correlated components across the different EEG recordings correspond to stimulus-following/-related neural responses. This assumption also delineates the context in which we propose SI-GCCA, i.e., all EEG recordings are recorded using the same natural stimulus and are synchronized in time. In that context, the assumption holds by design.

### v3

@@ 1 Introduction
In the past few years, we have seen a surge of BCI applications that are passive, i.e., tapping into the natural behavior of a user, single-trial, i.e., not requiring repetitions of the stimulus, and operate on natural sensory stimuli, such as speech/music and natural video [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. These new BCI paradigms can then be employed in much more mainstream application domains, such as hearing aids and consumer earphones [14, 3, 15, 4, 16], educational sciences [5, 11, 17], neuromarketing [6], or virtual reality environments [18]. Many of these applications involve decoding stimulus-following neural responses, for example, to quantify levels of absolute [8, 10, 19] or selective attention [14, 3] to a particular auditory or visual stimulus. The temporal dynamics of the stimulus then result in a so-called stimulus-following neural response (i.e., the neural tracking phenomenon [20]), which can be decoded from different neurorecording modalities such as electroencephalography (EEG) [20], magnetoencephalography (MEG) [21], or electrocorticography (ECoG) [22]. The EEG modality is particularly interesting because it is non-invasive, low cost and highly mobile [1].
However, decoding neural responses to natural stimuli is much more challenging from a signal processing perspective, given that they are much more unpredictable (resulting from the nature of the stimulus) and suffer from a very low SNR, as the targeted stimulus-following neural responses are buried under all kinds of non-neural and neural noise. Given the single-trial paradigm, averaging the responses across multiple trials is not an option anymore to deal with this extremely low SNR. Therefore, much more advanced, data-driven signal processing algorithms are required and being developed to enhance the targeted stimulus-following activity and suppress all other noise [23, 3, 24]. In this paper, we focus on the group analysis of such stimulus-following responses, i.e., we assume a set of synchronized neural responses to the same natural stimulus is available, e.g., from a group of subjects all attending to the same stimulus. This synchronization hypothesis roots in the previously mentioned neural tracking phenomenon of natural stimuli, where the coherent neural responses across subjects arise from the synchronized stimulus-following neural responses. Importantly, this synchronization across subjects refers only to these stimulus-locked responses that are typically earlier responses, appearing in the post-stimulus range [20, 22, 21, 14]. More complex, longer-latency responses that are, for example, related to emotional arousal or cognition are unlikely to be synchronous across subjects. This hypothesis of synchronicity for group decoding has been confirmed in various papers in literature using data-driven decoding algorithms (e.g., [6, 7, 8, 5, 25, 26, 13]). The group analysis of the stimulus-following neural responses can be a goal in itself, e.g., to decode a notion of group attention to the stimulus, as employed in educational neuroscience [5] or neuromarketing [6]. Alternatively, the group information can be leveraged to assist the decoding of stimulus-following neural responses on an individual level, using it as, e.g., a preprocessing technique to a priori improve the SNR [25, 26].
@@ 3 SI-GCCA/SI-corrCA
We now propose the stimulus-informed GCCA (SI-GCCA) technique for the group analysis of stimulus-following neural responses, in which we include the stimulus as side information, as visualized in Figure 1. The implicit assumption in SI-GCCA is, therefore, that the correlated components across the different EEG recordings correspond to (early) stimulus-following/-related neural responses, as explained in the introduction. This assumption also delineates the context in which we propose SI-GCCA, i.e., all EEG recordings are recorded using the same natural stimulus and are synchronized in time.
@@ 5.1 Amount of training data
A first important observation from Figure 3a-i, and from Figure 3 in general is that significant ISCs can be obtained using data-driven group decoding algorithms. This shows that synchronized stimulus-following responses across subjects exist and can be decoded. The latter has already been established in other studies on neural responses to natural stimuli [6, 8, 5], and the ISC values in Figure 3 are in line with those observed in other studies. In general, ISC and SC values are typically low because of the very low SNR of the stimulus-following neural responses w.r.t. the background EEG activity.